# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v3

@@ 6 Conclusion
Our code for the experiments with Diffusion Policy is available at https://github.com/YuejiangLIU/bid_diffusion and our code for the experiments with VQ-BET is available at https://github.com/Jubayer-Hamid/bid_lerobot. Additionally, detailed descriptions of our experimental setups are available in Appendix B, and complete proofs of our theoretical claims can be found in Appendix D.
@@ A.6 Choice of Distance Metrics
Appendix B Additional Details
B.1 Simulation Experiment Details
B.1.1 Environment Details
@@
B.2 Real-world Experiment Details
B.2.1 Dynamic Placing
@@ Results.
Appendix C Additional Discussions
@@
Appendix D Proofs
@@
where $\epsilon=\max_{X^{\prime}\neq X,X,Y}\{\mathcal{L}(f^{*}(X^{\prime}),S(X,Y)\}$ and $f^{*}=\arg\min_{f}\{\mathbb{E}_{X,Y}\left[\mathcal{L}(f(X),S(X,Y)\right]\}$
@@
D.1 Definitions
@@
Consider a $(c,h)$-policy (i.e., the policy has context length $c$ and action horizon $h$) whose probability of taking action $a_{t}$ at time $t$ in a chunk generated at $t$ is referred to as
$\pi_{(c,h)}:=\pi_{(c,h)}(a_{t}|s_{t-c:t}).$
On the other hand, consider a $(c+1,h)$-policy whose probability of taking action $a_{t}$ in a chunk generated at time $t$ is referred to as
$\pi_{(c+1,h)}:=\pi_{(c+1,h)}(a_{t}|s_{t-c-1:t}).$
Lastly, consider a $(k,1)$-expert whose probability of taking action $a_{t}$ at time $t$ is $\pi^{*}$.
@@
In particular, this is an equality when the state $s_{t-c-1}$, conditioned on $C$, takes on only one value almost surely.
@@
$\alpha_{b}\coloneqq\min_{\pi_{(c,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (8)
where $C=\{s_{t-h-c:t-h},a_{t-h:t-1}\}$ and $G=\{a_{t},s_{t-k:t-h-c-1},s_{t-h+1:t},z_{t-k:t}\}$. Clearly, $\alpha_{b}\geq 0$. In particular, $\alpha_{b}=0$ when $s_{t-h-d-c:t-h-c-1}$ can be deterministically inferred by $\pi_{(c,h)}$ or when the expert policy is independent of them.
Definition (Maximum Inference Disadvantage ($\epsilon_{f}$)). Consider the maximum divergence that can be accumulated by the ${(c,h+d)}$-policy from not knowing the recent states at time steps $s_{t-h-d+1:t-h}$ and let that be $\epsilon_{f}$. More formally, we say that, for fixed $C$ from Proposition 1, any states in $\mathcal{S}^{-}$, any $z_{t-k:t}$ and any $\hat{s}_{t-h-d+1:t-h}\neq{s}_{t-h-d+1:t-h}$:
@@
Here, $\pi_{(c+d,h)}\coloneqq\arg\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})|C]$ is the optimal $(c+d,h)$-policy.
To define $\alpha_{f}$ and $\epsilon_{b}$, we prove a second version of Proposition 6. Consider a $(c,h)$-policy whose probability of taking action $a_{t}$ at time $t$ in a chunk generated at $t$ is referred to as
$\pi_{(c,h)}:=\pi_{L(c,h)}(a_{t}|s_{t-c:t}).$
On the other hand, consider a $(c-1,h+1)$-policy whose probability of taking action $a_{t}$ in a chunk generated at time $t-1$ is referred to as
$\pi_{(c-1,h+1)}:=\pi_{(c-1,h+1)}(a_{t}|s_{t-c:t-1}).$
Lastly, consider a $(k,1)$-expert whose probability of taking action $a_{t}$ at time $t$ is $\pi^{*}$.
@@
$\min_{\pi_{(c,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}C\right]\leq\min_{\pi_{(c-1,h+1)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c-1,h+1)},\pi^{*})\Big{|}C\right]$
In particular, this is an equality when the state $s_{t}$, conditioned on $C$, takes on only one value almost surely.
@@
Definition (Expected Observation Advantage ($\alpha_{f}$)). Recall that we have two models: $\pi_{(c,h)}$ and $\pi_{(c,h+d)}$ and a hypothetical $(c,h+d)$-policy that has access to all the information both our learners have (as in Eq. 8 and Eq. 9). Observe that $\pi_{(c+d,h)}$ has access to more context than $\pi_{(c,h+d)}$, particularly the knowledge of states $s_{t-h-d+1:t-h}$. Therefore, we know, using Proposition 7, $\pi_{(c+d,h)}$ has lower divergence with respect to $\pi^{*}$ than $\pi_{(c,h+d)}$. We say that the advantage $\pi_{(c+d,h)}$ gets from the extra information is $\alpha_{f}$. More formally, we say that
$\alpha_{f}=\min_{\pi_{(c,h+d)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (10)
where $C=\{s_{t-h-d-c:t-h-d},a_{t-h-d:t-1}\}$ and $G=\{a_{t},s_{t-k:t-h-d-c-1},s_{t-h-d+1:t},z_{t-k:t}\}$. Clearly, $\alpha_{f}\geq 0$. In particular, $\alpha_{f}=0$ when $\pi_{(c,h+d)}$ can infer $s_{t-h-d+1:t-h}$ perfectly i.e. when the environment is completely static with $P_{f}=1$. This makes sense–in the static environment, observing these states does not provide any advantage since the optimal $\pi_{(c,h+d)}$ can infer these states anyway using the actions taken at those time steps.
@@
Here, $\pi_{(c+d,h)}\coloneqq\arg\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})|C]$ is the optimal $(c+d,h)$-policy.
The intuitive relationship between $\alpha_{f}$ and $\epsilon_{f}$ (and the same for $\alpha_{b}$ and $\epsilon_{b}$) holds:
@@
$\allowdisplaybreaks\alpha_{f}$ $=\min_{\pi_{(c,h+d)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$
$=\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P_{\text{correct inference}}\pi_{(c+d,h)}+P_{\text{incorrect inference}}{\pi}_{(c+d,h)},\pi^{*})\Big{|}C\right]$
$-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$
$\leq\min_{\pi_{(c+d,h)}}\{\mathbb{E}_{G}\left[P_{\text{incorrect inference}}\mathcal{L}({\pi}_{(c+d,h)}(\text{conditioning on incorrect inference}),\pi^{*})\Big{|}C\right]$
$+\mathbb{E}_{G}\left[P_{\text{correct inference}}\mathcal{L}(\pi_{(c+d,h)}(\text{conditioning on correct inference}),\pi^{*})\Big{|}C\right]\}$
$-\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (Convexity)
$\leq\mathbb{E}_{G}\left[P_{\text{incorrect inference}}\mathcal{L}(\hat{\pi}_{(c+d,h}^{*}(\text{conditioning on incorrect inference}),\pi^{*})\Big{|}C\right]$
$+\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)}^{*},\pi^{*})\Big{|}C\right]-\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)}^{*},\pi^{*})\Big{|}C\right]$ (Bounding probabilities by 1 and Lemma 5)
$\leq\mathbb{E}_{G}\left[P_{\text{incorrect inference}}\epsilon_{f}\mid C\right]$
@@
Here, $\pi_{(c+d,h)}^{*}\coloneqq\arg\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$. ∎
@@
Definition (Forward and Backward Inference). For a fixed time step $t$ and $C$ (as in Sec. 3.2), consider the time steps $\{t-h-d+1:t-h\}$. Define
$P_{f}(t^{\prime})\coloneqq P(S_{t^{\prime}}=g_{t^{\prime}}|S_{t^{\prime}-1}=g_{t^{\prime}-1},A_{t^{\prime}-1}=a_{t^{\prime}-1})$
@@
for all $t^{\prime}\in[t-h-d+1:t-h]$ with $g_{t^{\prime}},g_{t^{\prime}-1},a_{t^{\prime}-1}$ being the ground truth states and action in the deterministic environment. We assume that $P_{f}(t^{\prime})=1$ in a deterministic environment. In a stochastic environment, $P_{f}(t^{\prime})<1$ for all $t^{\prime}$ and as the stochasticity increases, these values decrease and approach 0. Then, we define
$P_{f}\coloneqq\sup\{P_{f}(t^{\prime})|t^{\prime}\in[t-h-d+1:t-h]\}.$
Now, consider the time steps $\{t-h-d-c:t-h-c-1\}$. Define
$P_{b}(t^{\prime})\coloneqq P(S_{t^{\prime}}=g_{t^{\prime}}|S_{t^{\prime}+1}=g_{t^{\prime}+1})$
for all $t^{\prime}\in[t-h-d-c:t-h-c-1]$. Since this is not conditioned on any action and as conditioning on $a_{t^{\prime}}$ reduces entropy, we assume that $P_{b}(t^{\prime})<1$ for all $t^{\prime}\in[t-h-d-c:t-h-c-1]$ in all environments. As stochasticity increases, $P_{b}(t^{\prime})$ decreases and approaches 0. Then, define
$P_{b}\coloneqq\sup\{P_{b}(t^{\prime})|t^{\prime}\in[t-h-d-c:t-h-c-1]\}.$
D.2 Discussion on Assumption 2
Before we prove the next result, we briefly discuss Assumption 2. Note that, using law of total probability, we can write:
$\pi_{(c,h)}(a_{t}|s_{t-h-c:t-h},a_{t-h:t-1})$
$=\sum_{\begin{subarray}{c}s_{t-k:t-h-c-1},\\ s_{t-h+1:t}\end{subarray}}\hat{P}(s_{t-k:t-h-c-1},s_{t-h+1:t}|s_{t-h-c:t-h},a_{t-h:t-1})\pi_{(k,1)}(a_{t}|s_{t-k:t},a_{t-h:t-1})$ (Law of Total Probability)
$=\mathbb{E}_{s_{t-k:t-h-c-1},s_{t-h+1:t}\sim\hat{P}}\left[\pi_{(k,1)}(a_{t}|s_{t-k:t})|s_{t-h-c:t-h},a_{t-h:t-1}\right].$
Here $\hat{P}$ is the policy’s learned (implicit) dynamics model. Assumption 2 allows us to replace $\hat{P}$ with the true dynamics model $P$. Given that we discuss optimal policies for each choice of action horizon trained with infinite data, this is a reasonable. Note that this does not make inference trivial - the policy learns the distribution but, given the distribution has non-zero entropy in a non-deterministic environment, the policy can still infer the wrong state.
Thus, using assumption 2, we can write the optimal policy as
$\pi_{(c,h)}=\mathbb{E}_{\hat{s}_{t-k:t-h-c-1},\hat{s}_{t-h+1:t}\sim P}\left[\pi_{(k,1)}|s_{t-h-c:t-h},a_{t-h:t-1}\right]$
where $P(S_{t+1}=s_{t+1}|S_{t}=s_{t},a_{t})$ is the environment’s transition dynamics and $\pi_{(k,1)}$ is the distribution of $a_{t}$ under a $(k,1)$-model.
D.3 Consistency-Reactivity Inequalities
Now we prove the Consistency-Reactivity Inequalities. We prove the upper and lower bound separately:

### v4

@@ 6 Conclusion
Our code for the experiments with Diffusion Policy is available at https://github.com/YuejiangLIU/bid_diffusion and our code for the experiments with VQ-BET is available at https://github.com/Jubayer-Hamid/bid_lerobot. Additionally, detailed descriptions of our experimental setups are available in Appendix C, and complete proofs of our theoretical claims can be found in Appendix E.
@@ A.6 Choice of Distance Metrics
Appendix B Additional Discussions
Appendix C Additional Details
C.1 Simulation Experiment Details
C.1.1 Environment Details
@@
C.2 Real-world Experiment Details
C.2.1 Dynamic Placing
@@ Results.
Appendix D Additional Discussions
@@
Appendix E Proofs
In this section, we will clearly write out the transition dynamics in the training environment to be $P_{\mathrm{train}}(s_{t+1}\mid s_{t},a_{t})$, the transition dynamics in the test environment to be $P_{\mathrm{test}}(s_{t+1}\mid s_{t},a_{t})$ and the implicit dynamics model of the training environment to be $\hat{P}_{\mathrm{train}}(s_{t+1}\mid s_{t},a_{t})$. Note that the implicit dynamics model attempts to learn the dynamics of the training environment and not the test environment since our policy is assumed to have been trained only on data from the training environment.
@@ Appendix E Proofs
where $\epsilon=\max_{X^{\prime}\neq X,X,Y}\{\mathcal{L}(f^{*}(X^{\prime}),S(X,Y)\}$ and $f^{*}=\arg\min_{f}\{\mathbb{E}_{X,Y}\left[\mathcal{L}(f(X),S(X,Y)\right]\}$.
@@
E.1 Assumptions
This assumption allows us to focus on the problem that is relevant to us and allows us to ignore edge cases. However, our analysis can be extended to the case where this assumption does not hold. In the case where $c+h\geq k$, the larger action chunk model will not get any advantage (making $\alpha_{b}=0$) since the additional states it has observed in time steps $\tau_{b}$ are irrelevant to $a_{t}$. Then, the longer horizon policy either attains the same expected loss the shorter horizon policy or suffer greater loss from not having observed the recent past states.
Assumption 2. An optimal $\pi_{c,h}$ must infer the unobserved states based on the observed states and actions by modeling the transition dynamics $\hat{P}_{\mathrm{train}}(s_{t^{\prime}}\mid s_{t^{\prime}-1},a_{t^{\prime}-1})$ accurately for all time step $t^{\prime}$. We consider this implicit model to be the same for both $\pi_{(c,h)}$ and $\pi_{(c,h+d)}$.
In other words, we can write using the law of total probability:
$\pi_{(c,h)}(a_{t}|s_{t-h-c:t-h},a_{t-h:t-1})$
$=\mathbb{E}_{s_{t-k:t-h-c-1},s_{t-h+1:t}\sim\hat{P}_{\mathrm{train}}}\left[\pi_{(k,1)}(a_{t}|s_{t-k:t})|s_{t-h-c:t-h},a_{t-h:t-1}\right]$
and
$\pi_{(c,h+d)}(a_{t}|s_{t-h-d-c:t-h-d},a_{t-h-d:t-h-1})$
$=\mathbb{E}_{s_{t-k:t-h-d-c-1},s_{t-h-d+1:t}\sim\hat{P}_{\mathrm{train}}}\left[\pi_{(k,1)}(a_{t}|s_{t-k:t})|s_{t-h-d-c:t-h-d},a_{t-h-d:t-1}\right].$
E.2 Definitions
@@
Consider a $(c,h)$-policy (i.e., the policy has context length $c$ and action horizon $h$) whose probability of taking action $a_{t}$ at time $t$ in a chunk generated at $t$ is referred to as $\pi_{(c,h)}:=\pi_{(c,h)}(a_{t}|s_{t-c:t}).$ On the other hand, consider a $(c+1,h)$-policy whose probability of taking action $a_{t}$ in a chunk generated at time $t$ is referred to as $\pi_{(c+1,h)}:=\pi_{(c+1,h)}(a_{t}|s_{t-c-1:t}).$ Lastly, consider a $(k,1)$-expert whose probability of taking action $a_{t}$ at time $t$ is $\pi^{*}$.
@@
In particular, this is an equality if and only if $s_{t-c-1}\sim P_{\mathrm{test}}(\cdot\mid C)$ and $\hat{s}_{t-c-1}\sim\hat{P}_{\mathrm{train}}(\cdot\mid C)$ take on only one and the same value almost surely.
@@
$\alpha_{b}\coloneqq\min_{\pi_{(c,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (8)
where $C$ is defined as in Proposition 1. Clearly, $\alpha_{b}\geq 0$. In particular, $\alpha_{b}=0$ when $s_{t-h-d-c:t-h-c-1}$ can be deterministically inferred by $\pi_{(c,h)}$ or when the expert policy is independent of them.
Definition (Maximum Inference Disadvantage ($\epsilon_{f}$)). Consider the maximum divergence that can be accumulated by the ${(c,h+d)}$-policy from not knowing the recent states at time steps $s_{t-h-d+1:t-h}$, and let that be $\epsilon_{f}$. More formally, we say that, for fixed $C$ from Proposition 1, any states in $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$, any $z_{t-k:t}$, and any $\hat{s}_{t-h-d+1:t-h}\neq{s}_{t-h-d+1:t-h}$, the following holds:
@@
Here, $\pi_{(c+d,h)}\coloneqq\arg\min_{\pi_{(c+d,h)}}\mathbb{E}[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})|C]$ is the optimal $(c+d,h)$-policy.
To define $\alpha_{f}$ and $\epsilon_{b}$, we prove a second version of Proposition 6. Consider a $(c,h)$-policy whose probability of taking action $a_{t}$ at time $t$ in a chunk generated at $t$ is referred to as $\pi_{(c,h)}:=\pi_{(c,h)}(a_{t}|s_{t-c:t}).$ On the other hand, consider a $(c-1,h+1)$-policy whose probability of taking action $a_{t}$ in a chunk generated at time $t-1$ is referred to as $\pi_{(c-1,h+1)}:=\pi_{(c-1,h+1)}(a_{t}|s_{t-c:t-1}).$ Lastly, consider a $(k,1)$-expert whose probability of taking action $a_{t}$ at time $t$ is $\pi^{*}$.
@@
$\min_{\pi_{(c,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c,h)},\pi^{*})\Big{|}C\right]\leq\min_{\pi_{(c-1,h+1)}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{(c-1,h+1)},\pi^{*})\Big{|}C\right].$
In particular, this is an equality if and only if the state $s_{t}\sim P_{\mathrm{test}}(\cdot\mid C)$ and $\hat{s}_{t}\sim\hat{P}_{\mathrm{train}}(\cdot\mid C)$ takes on only one and the same value almost surely.
@@
Definition (Expected Observation Advantage ($\alpha_{f}$)). Recall that we have two models: $\pi_{(c,h)}$ and $\pi_{(c,h+d)}$ and a hypothetical $(c+d,h)$-policy that has access to all the information both our learners have (as in Eq. 8 and Eq. 9). Observe that $\pi_{(c+d,h)}$ has access to more context than $\pi_{(c,h+d)}$, particularly the knowledge of states $s_{t-h-d+1:t-h}$. Therefore, we know, using Proposition 7, $\pi_{(c+d,h)}$ has lower divergence with respect to $\pi^{*}$ than $\pi_{(c,h+d)}$. We say that the advantage $\pi_{(c+d,h)}$ gets from the extra information is $\alpha_{f}$. More formally, we say that
$\alpha_{f}=\min_{\pi_{(c,h+d)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (10)
where $C$ is defined as in Proposition 1. Clearly, $\alpha_{f}\geq 0$. In particular, $\alpha_{f}=0$ when $\pi_{(c,h+d)}$ can infer $s_{t-h-d+1:t-h}$ perfectly. This makes sense–in the static environment, observing these states does not provide any advantage since the optimal $\pi_{(c,h+d)}$ can infer these states anyway using the actions taken at those time steps (assuming that the implicit dynamics model is accurate).
@@
Here, $\pi_{(c+d,h)}\coloneqq\arg\min_{\pi_{(c+d,h)}}\mathbb{E}[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})|C]$ is the optimal $(c+d,h)$-policy.
As a warmup, we see that the intuitive relationship between $\alpha_{f}$ and $\epsilon_{f}$ (and the same for $\alpha_{b}$ and $\epsilon_{b}$) holds:
@@
$\allowdisplaybreaks\alpha_{f}$ $=\min_{\pi_{(c,h+d)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c,h+d)},\pi^{*})\Big{|}C\right]-\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$
$=\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(P_{\text{correctinference}}\pi_{(c+d,h)}+P_{\text{incorrect inference}}{\pi}_{(c+d,h)},\pi^{*})\Big{|}C\right]$
$-\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$
$\leq\min_{\pi_{(c+d,h)}}\{\mathbb{E}\left[P_{\text{incorrect inference}}\mathcal{L}({\pi}_{(c+d,h)}(\text{conditioning on incorrect inference}),\pi^{*})\Big{|}C\right]$
$+\mathbb{E}\left[P_{\text{correct inference}}\mathcal{L}(\pi_{(c+d,h)}(\text{conditioning on correct inference}),\pi^{*})\Big{|}C\right]\}$
$-\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$ (Convexity)
$\leq\mathbb{E}\left[P_{\text{incorrect inference}}\mathcal{L}(\hat{\pi}_{(c+d,h}^{*}(\text{conditioning on incorrect inference}),\pi^{*})\Big{|}C\right]$
$+\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}^{*},\pi^{*})\Big{|}C\right]-\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}^{*},\pi^{*})\Big{|}C\right]$ (Bounding probabilities by 1 and Lemma 5)
$\leq\mathbb{E}\left[P_{\text{incorrect inference}}\epsilon_{f}\mid C\right]$
@@
Here, $\pi_{(c+d,h)}^{*}\coloneqq\arg\min_{\pi_{(c+d,h)}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right]$. ∎
We will provide a tighter bound after the proof of our main theoretical results.
@@
Definition (Forward and Backward Inference). For a fixed time step $t$ and $C$ (as in Sec. 3.2), consider the time steps $\tau_{f}\coloneqq\{t-h-d+1:t-h\}$ and $\tau_{b}\coloneqq\{t-h-d-c:t-h-c-1\}$.
Define
$P_{f}(t^{\prime})\coloneqq P_{\text{test}}(S_{t^{\prime}}=g_{t^{\prime}}\mid S_{t^{\prime}-1}=g_{t^{\prime}-1},A_{t^{\prime}-1}=a_{t^{\prime}-1})$
@@
for any $t^{\prime}\in\tau_{f}$ with $g_{t^{\prime}},g_{t^{\prime}-1},a_{t^{\prime}-1}$ being the ground truth states and action in the deterministic test environment. As such, $P_{f}(t^{\prime})=1$ in a deterministic environment. In a stochastic environment, $P_{f}(t^{\prime})<1$ for all $t^{\prime}$ and as the stochasticity increases, these values decrease and approach 0. Let
$\delta_{f}(t^{\prime})=\hat{P}_{\mathrm{train}}(g_{t^{\prime}}|g_{t^{\prime}-1},a_{t^{\prime}-1})-P_{f}(t^{\prime})$
where $g_{t^{\prime}},g_{t^{\prime}-1},a_{t^{\prime}-1}$ are still the ground truth in the deterministic test environment. Define $\hat{P}_{f}(t^{\prime})=P_{f}(t^{\prime})+\delta_{f}(t^{\prime})$. Similarly, define
$P_{b}(t^{\prime})\coloneqq P_{\text{test}}(S_{t^{\prime}}=g_{t^{\prime}}\mid S_{t^{\prime}+1}=g_{t^{\prime}+1})$
for any $t^{\prime}\in\tau_{b}$ where $g_{t^{\prime}},g_{t^{\prime}-1}$ are the ground truth states in the deterministic test environment. Then let
$\delta_{b}(t^{\prime})=\hat{P}_{\mathrm{train}}(g_{t^{\prime}}|g_{t^{\prime}+1})-P_{b}(t^{\prime})$
for any $t^{\prime}\in\tau_{b}$. Define $\hat{P}_{b}(t^{\prime})=P_{b}(t^{\prime})+\delta_{b}(t^{\prime})$.
Intuitively, $\delta_{f}(t^{\prime})$ and $\delta_{b}(t^{\prime})$ characterize the error that the implicit dynamics model has in capturing the test environment’s dynamics.
E.3 Consistency-Reactivity Inequalities
Now we prove the Consistency-Reactivity Inequalities.